OpenStudy (anonymous):

when the sun directly overhead, a hawk dives toward the ground with costan velocity of 5.30 m/s at 52.5 degree below the horizontal. calculate the speed of her shado on level ground

OpenStudy (anonymous):

We need to determine the horizontal component of velocity. \[\cos(\theta) = {v_h \over v}\]where \(\theta\) is the angle to the horizontal, \(v_h\) is the horizontal component of velocity, and \(v\) is the given velocity of the hawk.